# LNS2+RL: Combining Multi-agent Reinforcement Learning with Large Neighborhood Search in Multi-agent Path Finding

### v1

@@ 4.2 RL environment setup
We initially establish a MAPF task comprising $m$ agents, assigning each a unique start and a corresponding goal position. Using PP+SIPPS to compute the initial path set $P$ for the $m$ agents. Subsequently, a subset of agents is selected, and during the training phase of the MARL model, we keep the subset size $n$ (i.e., the neighborhood size) to 8. The PMDO re-planning task is then formulated by designating the chosen subset as the controlled agents of the MARL model, while the remaining agents act as soft obstacles. A training episode terminates either when all agents are on goal at the end of a timestep (the timestep is higher than the maximum length of the soft obstacle trajectories) or when the episode reaches the pre-defined timestep limit.
4.3 Observation and reward
Aligned with previous works Sartoretti et al. (2019), the environment is partially observable, restricting each agent to only access information within its self-centered field of view (FoV). This assumption enables the trained policy to scale to arbitrary world sizes while fixing the input dimensions of the neural network. The FoV contains global spatial and temporal information derived from the complete map structure and past/predicted trajectories of all soft obstacles and other agents.
Each agent’s current observation at timestep $t$ is divided into two parts. Here, we take agent $i$ as an example. The first part of the observation consists of thirty-one 2D matrices. Following methodologies from previous research Wang et al. (2023), where four matrices represent the actual/projected positions of hard obstacles, other agents, observable agents’ goals, and agent $i$’s own goal. Additionally, there are four binary matrices representing the shortest paths of agent $i$, calculated by the Dijkstra algorithm without considering other agents and soft obstacles. Another map that helps guide agent $i$ closer to its goal is the SIPPS map, which represents all vertices traversed by agent $i$’s SIPPS path $\hat{p}_{i}$ (computed only at the beginning of the re-planning task) during $t-15$ to $t+15$. To motivate MARL agents to refer to the SIPPS paths while allowing necessary detours to avoid collisions, we introduce an off-route penalty $ro$, that is
$\small ro_{t}^{i}=-\alpha\min\left\|v_{t}^{i}-v\right\|_{2},{v\in\left\{\left(\hat{p}_{i}\right)_{j}\mid t-15\leq j<t+15\right\}},$ (1)
where $\alpha$ is a predefined hyperparameter, $v_{t}^{i}$ is the position of agent $i$ at timestep $t$, $v$ represents all vertices contained in the SIPPS map of agent $i$.
To avoid collisions with hard and soft obstacles, we further introduce an occupancy duration map, a blank duration map, future trajectory maps, and predicted trajectory maps. The occupancy duration map denotes the ratio of time a vertex will be occupied by obstacles to the predefined episode length, while the blank duration map reflects the ratio of time a vertex will remain obstacle-free. The future trajectory maps, comprising nine matrices, represent the positions of soft obstacles from $t$ to $t+8$. The future trajectories of other agents are crucial to enhancing cooperation, but these trajectories are unknown in advance. Hence, we employ a straightforward yet efficient method for forecasting them. To predict the future trajectory of agent $j$, we first compute the spatial Euclidean distance between $v_{t}^{j}$ and all vertices on $\hat{p}_{j}$, and then determine the temporal distance as the absolute difference between $t$ and the timesteps of vertices on $\hat{p}_{j}$. The predicted position of agent $j$ at $t+1$ is then derived from a weighted sum of these spatial and temporal distances, with coefficients of 0.9 for spatial and 0.1 for temporal distances. If the vertex with the minimal sum distance is reachable in one action from $v_{t}^{j}$, it is considered the predicted position for $t+1$; otherwise, agent $j$ is assumed to remain stationary, with its actual position $v_{t}^{j}$ as the predicted position for $t+1$. The subsequent four vertices on the SIPPS path of the vertex with the minimal sum distance are designated as predicted positions from $t+2$ to $t+5$.
In path planning tasks involving a large number of agents, congestion often arises in bottleneck areas of the map, where all agents attempt to traverse the same set of vertices, leading to under-utilization of other vertices and significantly increased collision risks. To alleviate congestion, we introduce the vertex utilization map and the edge utilization maps, which visualize congestion and assist agents in identifying and avoiding high-density areas, promoting more evenly dispersed paths. Specifically, denote $U(v,t-\beta_{l},t+\beta_{h})$ and $U((v_{1},v_{2}),t-\beta_{l},t+\beta_{h})$ as the usage of vertex $v$ and edge $(v_{1},v_{2})$ during the time period from $t-\beta_{l}$ to $t+\beta_{h}$ , respectively. They are the sum of the number of times the soft obstacles traverse through the vertex/edge from $t-\beta_{l}$ to $t+\beta_{h}$, and the number of times agents traverse through the vertex/edge from $t-\beta_{l}$ to $t$. In our experiments, we set $\beta_{l}$ to 2 and $\beta_{h}$ to 15. The normalized value $\delta\frac{U(v,t-\beta_{l},t+\beta_{h})}{m}$ and $\delta\frac{U((v_{1},v_{2}),t-\beta_{l},t+\beta_{h})}{m}$ are then used as the value of each grid in the vertex utilization map and four edge utilization maps ($\delta=10$). Moreover, we further encourage agents to avoid these congested areas by giving them a small penalty $rc$ calculated by
$\small rc_{t}^{i}=\delta(\delta_{v}\frac{U(v,t-\beta_{l},t+\beta_{h})}{m}+\delta_{e}\frac{U((v_{1},v_{2}),t-\beta_{l},t+\beta_{h})}{m}).$ (2)
Here $\delta_{v}=0.225$, $\delta_{e}=0.075$ are used to balance between vertex and edge information.
The second part of the observation is an eight-element vector consisting of the normalized distance between $v_{t}^{i}$ and $g_{i}$ along the x-axis and y-axis, as well as the total Euclidean distance, the ratio of the number of CP in path planned by the MARL-planner compared to these in $\hat{p_{i}}$, the ratio of $t$ to the predefined episode length, the ratio of $t$ to the maximum length of $\hat{P}$, the ratio of agents who have reached goal, and agent $i$’s action at $t-1$.
Our reward structure is shown in Table 1. To expedite task completion, agents currently off-goal are penalized at each step, with its magnitude scaling up according to the task’s difficulty within the curriculum learning framework (refer Section 4.5 for details). Agents receive an extra penalty if the current timestep exceeds the maximum length of $\hat{P}$. Additionally, to foster exploration and learn effective policies, agents are penalized for revisiting their previous timestep’s location. Furthermore, we employ reward shaping based on the distance from the goal to guide learning while preserving the invariance of the optimal policy Ng et al. (1999).
Table 1: Reward structure
@@ 4.3 Observation and reward
Move/Stay
(not on goal)
Stay
(on goal)
Exceed the
maximum length of $\hat{P}$
Return
Collision
with obstacles
Avoid
congestion
@@
Reward [-0.4,-0.5,-0.6] 0.0 -0.2 -0.4 -1.5 detailed in equation 2 detailed in equation 1
4.4 Network structure
Figure 2: Network structure. Green, purple, and red modules represent the memory encoder, observation encoder, and output heads, respectively. $h_{t-1}^{i}$ represents the hidden state output by the LSTM unit at the previous time step.
@@ 4.4 Network structure
Figure 2 illustrates the network architecture of our MARL model, encompassing three primary components: memory encoder, observation encoder, and output heads. Given our assumption of a homogeneous team, we rely on weight sharing across all agents, enabling the model to scale effectively with the number of agents during evaluation.
@@
Understanding the underlying dynamics of the environment is crucial in PMDO tasks. Therefore, in addition to utilizing occupied duration map, blank duration map, future trajectory maps, and predicted trajectory maps to capture future temporal information, we incorporate observations from the past three time steps as input to the neural network to provide a more comprehensive understanding of how the environment changes over time. The aggregated observation first passes through a Convolutional Long Short-Term Memory (ConvLSTM) Shi et al. (2015) unit capable of capturing spatial and temporal relationships and then concatenates (Cat) its last hidden state with the current observation $o_{t}^{i}$. The concatenated data then goes through an observation encoder with two ResNet blocks He et al. (2016), followed by an average pooling and Layer Normalization (LN) to produce a one-dimensional feature vector. In parallel, the eight-element vector $v_{t}^{i}$ is processed through a Fully Connected (FC) layer, accompanied by two LN. The processed inputs are then concatenated and passed through two FC layers equipped with a residual shortcut, followed by a LN, and ultimately fed into a Long Short-Term Memory (LSTM) cell. The LSTM’s output is further processed by a FC layer, producing a vector $\dot{o}_{t}^{i}$ for imitation learning. Subsequently, $\dot{o}_{t}^{i}$ passes through a softmax activation to derive the agent’s policy and a sigmoid activation to yield a vector $\ddot{o}_{t}^{i}$ for supervised invalid action loss. Another FC layer also utilizes the LSTM’s output as input to generate the predicted state value.
4.5 Training
Our MARL model is trained by Proximal Policy Optimization (PPO) Schulman et al. (2017). The total loss of the MARL model is $L_{final}=\omega_{v}\cdot L_{v}+\omega_{\pi}\cdot L_{\pi}+\omega_{invalid}\cdot L_{invalid}-\omega_{e}\cdot L_{e}$, where $L_{v}$ and $L_{\pi}$ are the standard critic and actor loss in PPO. $L_{e}$ is the entropy of the agent’s policy, suppressing its decline shown to encourage exploration and inhibit premature convergence. $L_{invalid}$ is a binary cross-entropy loss (supervised invalid action loss) designed to reduce the logarithmic likelihood of selecting an invalid action. $\omega_{v}$, $\omega_{\pi}$, $\omega_{invalid}$, and $\omega_{e}$ are all manually set hyperparameters.
The training of the MARL model is divided into two stages. As described in Section 4.2, both stages commence by generating a high-level MAPF task and obtaining the initial path set $P$ through PP+SIPPS. Subsequently, a series of PMDO re-planning tasks are generated in different manners and used to train the MARL model. To enhance the agents’ generalization ability, we expose them to diverse scenarios during training by uniformly sampling the world size, hard obstacle rate, and the total number of agents for the MAPF task from a variety of values.
In the first training stage, when constructing each PMDO task, the MARL-controlled agents are chosen randomly, bypassing ALNS. After an MARL episode ends, the path set $P$ is updated with the SIPPS paths $\hat{P}$ calculated at the beginning of the episode, regardless of whether these new paths reduce CP in $P$. PMDO re-planning tasks are continually generated and only switch to the next new high-level MAPF task when the the maximum iteration limit is reached. Curriculum learning Bengio et al. (2009) is adopted during the first stage of training, allowing agents to rapidly acquire fundamental skills from simpler tasks before progressively facing more challenging scenarios. Specifically, we increase the task difficulty by increasing the hard obstacles rate and the total number of agents in the high-level MAPF tasks and advancing to greater difficulties once the predetermined volume of training data is accumulated. As task difficulty increases, the advantage of SIPPS weakens, and the agents’ preference for safely wandering in a certain area increases. Therefore, we gradually lower the value of $\alpha$ in the off-route penalty to reduce references to SIPPS paths and gradually increase the action cost to incentivize agents to move toward their goal. Additionally, we employ imitation learning (behavior cloning based on the centralized planner ODrM*) at the beginning of each task difficulty in the first stage to explore high-quality regions of the agents’ state-action space.
Upon reaching the manually predefined number of training steps, the training transitions to the second stage. This stage aims to acclimate the MARL policy to scenarios it will face when inserted into the LNS2+RL algorithm. Therefore, both the selection of MARL-controlled agents and the method for updating path set $P$ are exactly the same as the LNS2+RL algorithm’s execution (refer to Section 5 for specifics). A new high-level MAPF task is generated either when $P$ achieves a collision-free state or upon hitting the maximum iteration limit.
5 LNS2+RL
Algorithm 1 LNS2+RL
  Input: A MAPF instance $I$
  $P=\left\{p_{i}\mid a_{i}\in A\right\}\leftarrow$ runlnitialSolver$(I)$
  $\forall w_{h}\in W,w_{h}=1$; build collision graph $G_{c}=(V_{c},E_{c})$; $Q=\emptyset$ and $|Q|\leq\mu$
  $t_{m}=\max\{|p_{i}|:p_{i}\in P\}$, $t_{l}=d_{l}\cdot t_{m}$, $t_{h}=d_{h}\cdot t_{m}$
  while runtime limit not exceeded and $CP(P)>0$ do
     $h\leftarrow$ selectDestroyHeuristic$(W)$
     $A_{s},P^{-}\leftarrow$ selectAgentSet$(h,I,P,G_{c})$
     if $|Q|<\mu$ or $\bar{Q}>\rho$ then
        $P^{rl},\hat{P}\leftarrow$ runMARLPlaner$\left(I,P\backslash P^{-},A_{s},t_{l}\right)$; $A_{r}\leftarrow\{a_{i}\in A_{s}:p_{i,\text{end}}^{rl}\neq g_{i}\}$
        $P^{rl}\leftarrow$ runPriorityPlaner$\left(I,P\backslash P^{-},P^{rl},A_{r},t_{l}\right)$; $A_{h}\leftarrow\{a_{i}\in A_{s}:|p_{i}^{rl}|\geq t_{h}\}$
        $P^{rl}\leftarrow$ runPriorityPlaner$\left(I,P\backslash P^{-},P^{rl},A_{h},0\right)$
        Update $Q$ accoridng to if $CP(P^{rl})$geq $CP(\hat{P})$
        if $CP(\left(P\backslash P^{-}\right)\cup P^{rl})<=CP(P)$ then
           $P\leftarrow\left(P\backslash P^{-}\right)\cup P^{rl}$
           Update the collision graph $G_{c}$
           Update $w_{h}$
           continue
        end if
     else
        $\hat{P}\leftarrow$ runPriorityPlaner$\left(I,P\backslash P^{-},\emptyset,A_{s},0\right)$
     end if
     if $CP(\left(P\backslash P^{-}\right)\cup\hat{P})<=CP(P)$ then
        $P\leftarrow\left(P\backslash P^{-}\right)\cup\hat{P}$
        Update the collision graph $G_{c}$
     end if
     Update $w_{h}$
  end while
  return $P$
The strengths and weaknesses of PP+SIPPS and the MARL-based planner are distinct yet complementary. PP+SIPPS is fast and effective for simple tasks but tends to generate paths with more soft collisions in complex tasks. Conversely, our MARL-based planner, although slower, excels at planning paths with fewer collisions in difficult tasks. Therefore, our LNS2+RL algorithm retains the overarching framework of LNS2 but uses the MARL-based planner in the initial, more challenging stages of iterative planning for better solution quality (fewer collisions). It then adaptively switches to PP+SIPPS to quickly clean up the remaining collisions, thus balancing solution quality with the speed of the re-planning task solvers. Moreover, the key limitation in the one-shot SR of MARL-based planners is that while most agents have found high-quality paths, a few fail to reach their goals within the constrained episode length. To address this, we employ PP+SIPPS to supplement paths for these few agents, ensuring the high quality and completeness of the final solution.
The pseudocode for the LNS2+RL algorithm is presented in Algorithm 1. Specifically, for solving a MAPF instance $I$, LNS2+RL starts by invoking the initial solver (PP+SIPPS) to generate a path set $P$ containing collisions (lines 2), initializing the weights $W$ for the three neighborhood selection methods, and constructing the collision graph $G_{c}$ (lines 3). Furthermore, LNS2+RL introduces a queue $Q$ with a maximum capacity of $\mu$ to store the relative solution quality from PP+SIPPS and the MARL-planner (line 3). Subsequently, LNS2+RL counts the maximum length of individual paths in the initial solution ($t_{m}$), sets the timestep threshold $t_{l}$ for pausing the MARL-planner, along with the maximum length $t_{h}$ of each hybrid path, where the constant value $d_{h}>d_{l}$ (line 4). If the termination criteria are not met, LNS2+RL iteratively selects a subset of agents $A_{s}\subseteq A$ using the ALNS, and removes their paths $P^{-}=\left\{p_{i}\in P:a_{i}\in A_{s}\right\}$ from $P$, treating the remaining part of the path set $P\backslash P^{-}$ as fixed to construct a PMDO re-planning task (lines 6-7). The neighborhood size (the size of $A_{s}$) is fixed at 8 (equivalent to the number of agents used in MARL training) or 16. If the size of $Q$ is below than $\mu$, or its average exceeds another threshold $\rho$, LNS2+RL infers that the solution quality of the MARL-planner significantly exceeds that of PP+SIPPS at this stage (line 8). Therefore, LNS2+RL chooses the trained MARL-planner as the re-planning solver. The function in line 9 uses the MARL-planner to predict actions until timestep $t_{l}$ (or until the re-planning task is fully resolved, i.e., all agents have reached their goal without future soft collisions). It yields two path sets: $\hat{P}$, derived by PP+SIPPS at the onset of the re-planning task, acting as the reference paths for the MARL policy, and path set $P^{rl}$, computed by the MARL policy. For incomplete paths in $P^{rl}$ where agents haven’t reached their goals, we employ PP+SIPPS to plan the remaining paths from $t_{l}$ and merge these new paths with the existing ones (Line 10). Subsequently, all path lengths in $P^{rl}$ are reviewed, and those exceeding $t_{h}$ are directly replaced with paths re-planned by PP+SIPPS from timestep zero (lines 11). This operation aims to remove non-contributory segments planned by the MARL-planner, thus reducing the SoC of the final solution $P$ and planning time for subsequent iterations. Next, LNS2+RL updates $Q$ by appending 0 if $CP(P^{rl})\geq CP(\hat{P})$, and 1 otherwise (lines 12). Finally, LNS2+RL first checks whether $P^{rl}$ meets the update condition (line 13), namely if the CP of the repaired path set $\left(P\backslash P^{-}\right)\cup P^{rl}$ is no larger than that of the old path set. If not, LNS2+RL continues to check whether $\hat{P}$ meets the update condition (line 22). The MARL-planner continues to re-plan until either the switching condition or termination condition is met. Once the re-planning solver switches to PP+SIPPS, it cannot switch back to the MARL-planner and always addresses remaining collisions using the same procedure as LNS2.
6 Experiments
We compare LNS2+RL with vanilla LNS2 Li et al. (2022) as well as a representative set of scalable high-quality MAPF algorithms: the bounded-suboptimal algorithm EECBS Li et al. (2021) (with a suboptimality factor $w=5$), the complete but unbounded-suboptimal algorithm LACAM Okumura (2023), the MARL-based algorithm SCRIMP Wang et al. (2023), and the priority-based planner (PP) with random restarts (using SIPP Phillips & Likhachev (2011) to plan paths for single agents) Cohen et al. (2018). Here, we only show the results of LNS2, EECBS, and LACAM because the SR of SCRIMP and PP are 0% in most tasks, and they cannot solve even the simplest tasks in the random-small map. Both LNS2+RL and LNS2 use PP+SIPPS to find the initial solution and ALNS to generate neighborhoods. Code for LNS2, EECBS, LACAM was written in C++, and LNS2+RL was written in both C++ and Python and connected using pybind (Our full training and testing code is available at https://github.com/marmotlab/LNS2-RL). All experiments were conducted on a server equipped with one Nvidia GeForce RTX 3090 GPU and an AMD Ryzen 9 7950X3D 16-Core Processor. To obtain results faster, LNS2+RL, LNS2, EECBS used Ray to test multiple MAPF tasks simultaneously. However, due to high memory demands (we limited the maximum memory to 100GB per task), LACAM was tested using a single process.
6.1 Results for multiple map types
@@
Figure 3: Result on different map types. The success rate is the percentage of tasks fully solved within the planning time constraints (i.e., where all agents reached their goal without collisions). The remaining collision pairs metric represents the number of CP remaining in the overall solution after reaching the time constraint and is only available for LNS2+RL and LNS2. The confidence interval (shaded area) represents one standard deviation of the remaining CP. All three random maps have a 17.5% obstacle density. Except for the random-small map with a size of $10\times 10$ and the random-large map with a size of $50\times 50$, all other five maps have a size of $25\times 25$. For maps with sizes $10\times 10$, $25\times 25$, and $50\times 50$, the time constraints for each task are 100 seconds, 600 seconds, and 5000 seconds, respectively, with test counts of 100, 50, and 20. The neighborhood size of LNS2+RL is 16 on random-large maps and 8 on all other maps.
The evaluation results presented in Figure 3 demonstrate that LNS2+RL consistently achieves equal or higher SR compared to LNS2 across all tasks, except for those on the empty map with 77.5% agent density. It is noteworthy that the improvement of LNS2+RL on CP is more pronounced. In tasks with 65% agent density on the random-small map, despite both algorithms achieving the same SR, LNS2+RL remarkably reduces the remaining CP by 26.3% compared to LNS2.
Table 2: Sum of costs on random-small map with 50%, 55%, 60%, 65% agent density (accounting only for successful tasks).

### v2

@@ Introduction
We initially establish a MAPF task comprising $m$ agents, and use PP+SIPPS to compute the initial path set $P$ for these agents. Subsequently, a subset of agents is selected, and during the training phase of the MARL model, we keep the subset size $n$ (i.e., the neighborhood size) to 8. The PMDO replanning task is then formulated by designating the chosen subset as the controlled agents of the MARL model while the remaining agents act as soft obstacles. Throughout the remainder of this section, we refer to the selected agents simply as agents and to the others as soft obstacles. A training episode terminates either when all agents are on goal at the end of a timestep (and the timestep is larger than the maximum length of the soft obstacle paths) or when the episode reaches the pre-defined timestep limit.
@@
Move/Stay (not on goal)
Stay (on goal)
Exceed max length of $\hat{P}$
Revisit
Collision with obstacles
Avoid congestion
@@
Reward [-0.4,-0.5,-0.6] 0.0 -0.2 -0.4 -1.5 Eq. 2 Eq.1
Table 1: Reward structure.
Observation and Reward
Aligned with previous works (Sartoretti et al. 2019), the environment is assumed to be partially observable for each agent, restricting it to only access information within its self-centered field of view (FoV). This assumption enables the trained policy to scale to arbitrary world sizes while fixing the input dimensions of the neural network.
Each agent’s observation is divided into two parts. Here, we take agent $i$ and timestep $t$ as an example. The first part of the observation is an eight-element vector consisting of the normalized distance between $c_{t}^{i}$ (the position of agent $i$ at timestep $t$) and $g_{i}$ along the x-axis and y-axis, as well as the total Euclidean distance, the ratio of the number of CP in the path planned so far by the MARL-planner compared to that in $\hat{p_{i}}$ (agent i’s SIPPS path), the ratio of $t$ to the predefined episode length, the ratio of $t$ to the maximum length of $\hat{P}$, the ratio of agents who have reached goal, and agent $i$’s action at $t-1$. The section part of the observation consists of thirty-one 2D matrices (Summary of these metrics is in the supplementary material). These matrices can be classified into four categories.
1. Design from prior work
Building on previous research (Wang et al. 2023), we introduce four matrices to represent the positions of hard obstacles, other agents, observable agents’ goals, and agent $i$’s own goal. Additionally, there are four binary matrices (one for each move action) that indicate whether selecting the action will bring agent $i$ closer to its goal, as determined by the backward Dijkstra algorithm.
2. Reference path
To better guide agent $i$ closer to its goal, we introduce the SIPPS map, which represents all cells traversed by agent $i$’s SIPPS path $\hat{p}_{i}$ (computed only at the beginning of the replanning task) during timesteps $t-15$ to $t+15$. To motivate agents to refer to the SIPPS paths while allowing necessary detours to avoid collisions, we define an off-route penalty $ro_{t}^{i}$:
$ro_{t}^{i}=-\alpha\min\left\|c_{t}^{i}-c\right\|_{2},{c\in\left\{\left(\hat{p}_{i}\right)_{j}\mid t-15\leq j<t+15\right\}},$ (1)
where $\alpha$ is a predefined hyperparameter, $c$ represents all cells contained in the SIPPS map of agent $i$.
3. Avoid collisions
To avoid collisions with hard and soft obstacles (with known paths), we further introduce future path maps of soft obstacles, and an occupancy duration map along with its complementary blank duration map. The future path maps, comprising nine matrices, represent the positions of soft obstacles from $t$ to $t+8$. The occupancy duration map denotes the ratio of time a cell will be occupied by obstacles to the predefined episode termination length, while the blank duration map reflects the ratio of time a cell will remain obstacle-free.
The future paths of other agents are crucial to enhancing cooperation, but these paths are unknown in advance. Hence, we introduce predicted path maps of other agents and employ a straightforward yet efficient method for forecasting them. To predict the future path of agent $j$, we first calculate the spatial Euclidean distance between $c_{t}^{j}$ and all cells on $\hat{p}_{j}$, and treat the absolute difference of their timestep as the temporal distance. Next, we compute a weighted sum of the spatial and temporal distances for each cell, with weights of 0.9 for spatial distance and 0.1 for temporal distance. If the cell with the smallest weighted sum can be reached from $c_{t}^{j}$ in a single step, it is considered the predicted position at $t+1$; otherwise, agent $j$ is assumed to remain stationary, and its current position $c_{t}^{j}$ is regarded as the predicted position at $t+1$. Subsequently, the next four cells on $\hat{p}_{j}$ starting from the cell with the smallest weighted sum are designated as the predicted positions from $t+2$ to $t+5$.
4. Alleviate congestion
In path planning tasks involving a large number of agents, congestion often arises in bottleneck areas of the map, where all agents attempt to traverse the same set of cells, leading to under-utilization of other cells and significantly increased collision risks. To alleviate congestion, we introduce the cell utilization map and the direction utilization maps, which visualize congestion and assist agents in identifying and avoiding high-density areas, promoting more evenly dispersed paths. Specifically, denote $U(c)$ and $U(c_{1},c_{2})$ as the usage of cell $c$ and direction $(c_{1},c_{2})$ during the time period from $t-2$ to $t+15$. They represent, respectively, the total timesteps during that time period in which soft obstacles and agents occupy cell $c$ and traverse direction $(c_{1},c_{2})$. The normalized value $10\frac{U(c)}{m}$ and $10\frac{U(c_{1},c_{2})}{m}$ are then used as the value of each grid in the cell utilization map and four direction utilization maps. Moreover, we further encourage agents to avoid these congested areas by giving them a small penalty $rc_{t}^{i}$ calculated by
$rc_{t}^{i}=10(\delta_{c}\frac{U(c)}{m}+\delta_{d}\frac{U(c_{1},c_{2})}{m}).$ (2)
Here $\delta_{c}=0.225$ and $\delta_{d}=0.075$ are used to balance between cell and direction information.
Our reward structure is shown in Table 1. To expedite task completion, agents currently off-goal are penalized at each step, with its magnitude scaling up according to the task’s difficulty within the curriculum learning framework (refer to “Training process“ subsection for details). Agents receive an extra penalty if the current timestep exceeds the maximum length of the current $\hat{P}$. Additionally, to foster exploration and learn effective policies, agents are penalized for revisiting their previous timestep’s location. Furthermore, we employ reward shaping based on the distance from the goal to guide learning while preserving the invariance of the optimal policy.
Network Structure
@@ 4. Alleviate congestion
Figure 1 illustrates the network architecture of our MARL model. Given our assumption of a homogeneous team, we rely on weight sharing across all agents, enabling the model to scale effectively with the number of agents during evaluation.
@@
Understanding the underlying dynamics of the environment is crucial in PMDO tasks. Therefore, we additionally incorporate observations from the past three time steps as input to provide a more comprehensive understanding of how the environment changes over time. The aggregated observation first passes through a Convolutional Long Short-Term Memory (ConvLSTM) (Shi et al. 2015) unit capable of capturing spatial and temporal relationships and then concatenates its last hidden state with the current observation $o_{t}^{i}$. The concatenated data and the eight-element vector $v_{t}^{i}$ are processed through multiple neural network layers to generate a vector $\dot{o}_{t}^{i}$ for imitation learning and a predicted state value. Subsequently, $\dot{o}_{t}^{i}$ passes through a softmax activation to derive the agent’s policy and a sigmoid activation to yield a vector $\ddot{o}_{t}^{i}$ for supervised invalid action loss.
Figure 1: Network structure. AP, Cat, LN, and FC represent average pooling, concatenation, layer normalization, and fully connected layer, respectively. $h_{t-1}^{i}$ represents the hidden state output by the LSTM unit at the previous time step.
@@
Figure 2: Result on different map types and visualization of representative maps. On the x-axis, the percentage represents agent density, while the number below it represents the number of agents. The success rate is the percentage of tasks that are fully solved before the termination condition is met. The remaining colliding pairs metric represents the number of CP remaining in the overall solution after reaching the time constraint and is only available for LNS2+RL and LNS2. The shaded area represents one standard deviation of the remaining CP. All three random maps have a 17.5% obstacle density. Except for the random-small map with a size of $10\times 10$ and the random-large map with a size of $50\times 50$, all other five maps have a size of $25\times 25$. For maps with sizes $10\times 10$, $25\times 25$, and $50\times 50$, the test counts are set to 100, 50, and 20, respectively, and the time constraints for each task, except when using SCRIMP, are set to 100 seconds, 600 seconds, and 5,000 seconds. To ensure consistency with the original paper, SCRIMP uses the maximum timestep as the termination condition, with maximum timesteps of 356, 556, and 1,024 for maps sized $10\times 10$, $25\times 25$, and $50\times 50$, respectively. Figure 2(h) shows the visualization of the random-medium (top left), maze (top right), room (bottom left), and warehouse (bottom right) map.
Training Process
Our MARL model is trained by Proximal Policy Optimization (PPO) (Schulman et al. 2017). The total loss of the MARL model is $L_{final}=\omega_{v}\cdot L_{v}+\omega_{\pi}\cdot L_{\pi}+\omega_{invalid}\cdot L_{invalid}-\omega_{e}\cdot L_{e}$, where $L_{v}$ and $L_{\pi}$ are the standard critic and actor loss in PPO. $L_{e}$ is the entropy of the agent’s policy, suppressing its decline shown to encourage exploration and inhibit premature convergence. $L_{invalid}$ is a binary cross-entropy loss (supervised invalid action loss) designed to reduce the logarithmic likelihood of selecting an invalid action. $\omega_{v}$, $\omega_{\pi}$, $\omega_{invalid}$, and $\omega_{e}$ are all manually set hyperparameters. Details regarding the hyperparameters, computational resources, and training duration are provided in the supplementary materials.
The training of the MARL model is divided into two stages. Both stages commence by generating a high-level MAPF task and obtaining the initial path set $P$ through PP+SIPPS. Subsequently, a series of PMDO replanning tasks are generated to train the MARL model. In the first training stage, to reduce training time, the MARL-controlled agents are randomly selected when constructing each PMDO task, bypassing adaptive neighborhood selection. After a MARL episode ends, the path set $P$ is updated with the SIPPS paths $\hat{P}$ calculated at the beginning of the episode, regardless of whether these new paths reduce CP in $P$. PMDO replanning tasks are continually generated and only switch to the next new MAPF task when the maximum iteration limit is reached. Curriculum learning is adopted during the first stage of training, allowing agents to rapidly acquire fundamental skills from simpler tasks before progressively facing more challenging scenarios. Specifically, we increase the task difficulty by increasing the hard obstacles rate and the total number of agents (MARL-controlled agents and soft obstacles) in the high-level MAPF tasks and advancing to greater difficulties once the predetermined volume of training data is accumulated. As task difficulty increases, the advantage of SIPPS weakens, and the agents’ preference for safely wandering in a certain area increases. Therefore, we gradually lower the value of $\alpha$ in the off-route penalty to reduce references to SIPPS paths and gradually increase the action cost to incentivize agents to move toward their goal. Additionally, we employ imitation learning (behavior cloning based on the centralized planner ODrM* (Wagner and Choset 2015)) at the beginning of each task difficulty in the first stage to explore high-quality regions of the agents’ state-action space. Upon reaching the manually predefined number of training steps, the training transitions to the second stage. This stage aims to acclimate the MARL policy to scenarios it will face when inserted into the LNS2+RL algorithm. Therefore, both the selection of agents and the method for updating path set $P$ are exactly the same as the LNS2+RL algorithm’s execution.
Experiments
We compare LNS2+RL with vanilla LNS2111Code from https://github.com/Jiaoyang-Li/MAPF-LNS2 as well as a representative set of MAPF algorithms: the bounded-suboptimal algorithm EECBS (with a suboptimality factor $w=5$) [^2] Code from https://github.com/Jiaoyang-Li/EECBS, the complete but unbounded-suboptimal algorithm LaCAM [^3] Code from https://github.com/Kei18/lacam, and the pure MARL-based algorithm SCRIMP [^4] Code from https://github.com/marmotlab/SCRIMP. Code for LNS2, EECBS, LaCAM was written in C++, SCRIMP was written in Python, and LNS2+RL was written in both C++ and Python and connected using pybind. To obtain results faster, LNS2+RL, LNS2, EECBS used Ray to test multiple MAPF tasks simultaneously. However, due to high memory demands (we limited the maximum memory to 100GB per task), LaCAM and SCRIMP were tested using a single process. The same MARL model was used in all evaluations, and it was trained on maps with random obstacles.
Comparison with Baselines
The evaluation results in Figure 2 show that LNS2+RL consistently achieves equal or higher SR compared to LNS2 across all tasks, except for those on the empty map with 77.5% agent density (484 agents). It is noteworthy that the improvement of LNS2+RL on CP is more pronounced. In tasks with 65% agent density (65 agents) on the random-small map, despite both algorithms achieving the same SR, LNS2+RL remarkably reduces the remaining CP by 26.3% compared to LNS2. The structural complexity of maps significantly influences algorithm performance. Despite being trained on random maps without prior exposure to environments with complex obstacle structures, such as mazes, rooms, and warehouses, LNS2+RL demonstrates remarkable improvements in SR on these three types of maps. Conversely, LaCAM excels on maps with simple or no obstacle structures, achieving a 100% SR on empty maps, but close to 0% SR on maps with complex obstacle structures. Due to the high difficulty of the selected tasks, the SR of EECBS and SCRIMP remains at 0% for most tasks. Furthermore, taking the tasks in Table 2 as an example, the SoC for LNS2+RL only averages 1.03% higher than that of LNS2. Despite LaCAM displaying higher SR on the random-small map, its SoC notably lags behind. Conversely, EECBS consistently achieves the best SoC but at the expense of obtaining the lower SR. SCRIMP exhibits the poorest performance in both SR and SoC. Moreover, LNS2+RL and LNS2 offer the added benefit of maintaining consistent memory usage, in contrast to the other three baselines, whose memory usage escalates rapidly over time due to the generation of longer paths, larger search frontiers, and more collisions.